# R-matrix valued Lax pair for elliptic Calogero-Inozemtsev system

### v2

@@ 3.1 The model and the Lax pair
Here we recall the Lax pair for the Calogero-Inozemtsev model (1.11) with five arbitrary constants $g,\nu_{0},\nu_{1},\nu_{2},\nu_{3}\in\mathbb{C}$ Equations of motion are of the form:
@@
In [13] $3n\times 3n$ Lax representation was suggested. Some particular cases (not arbitrary constants) were known previously from [25], see also [4]. Here we are going to deal with $2n\times 2n$ Lax representation proposed by K. Takasaki in [31]. See also [3], where this Lax pair arises in the non-relativistic limit, and we use the expression from [3]. The description of the spectral curve can be found in [1], and the case of $n=1$ was also studied in [33].
@@
where all entries $L^{ab}(z)$ and $M^{ab}(z)$, $a,b=1,2$ are $n\times n$ matrices (below $i,j=1,...,n$):
@@
where the functions $\phi$ and $f$ are given by (2.5) and (A.7) respectively and the following notations are used:
@@
$\begin{array}[]{c}\phi(z,u-w)(v^{\prime}(0,w)-v^{\prime}(0,u))=\\ \\ 2v(-z,w)f(z,u+w)+2v(z,u)f(-z,u+w)+v^{\prime}(-z,w)\phi(z,u+w)+v^{\prime}(z,u)\phi(-z,u+w)\,,\end{array}$ (3.23)
@@ 3.2 Lax equation
The explicit form can be obtained from (3.30) by replacing $q_{j},p_{j}$ to $-q_{j},-p_{j}$.
@@ 4.1 Associative Yang-Baxter equation and $K$-operators
Here the symmetric group $S_{N}$ acts by exchanging the variables $x_{i}$ and $\hat{s}_{ij}$ denotes the corresponding transposition
@@
Introduce another one operator $\tilde{K}$:
@@
[…]
Let us introduce notations $F_{ij}^{z}(q)$ (c.f. (2.8)) and $\tilde{F}_{ij}^{z}(q)$ for the corresponding partial derivatives of $R_{ij}^{z}(q)$ and $\tilde{R}_{ij}^{z}(q)$:
@@
[…]
@@ 4.2 ${\rm BC}_{n}$ associative Yang-Baxter equations with parameters
In order to define the associative classical Yang-Baxter algebra of type ${\rm A}_{n}$ generated by the elements $r_{ij}$ for $1\leq i\neq j\leq n$, we extend (4.29) with the antisymmetry condition and commutativity relation:
@@
The equation (1.8) with parameters was suggested by A. Polishchuk in [27]. The relation (4.29) looks as the half of the classical Yang-Baxter equation. From (4.29) and (4.30) the classical Yang-Baxter equation follows:
$[r_{12},r_{13}]+[r_{12},r_{23}]+[r_{13},r_{23}]=0.$ (4.31)
@@
[…]
@@
$\begin{array}[]{ccc}R_{ij}^{z}(q_{i}-q_{j})R_{jk}^{w}(q_{j}-q_{k})=\left(\phi(z,x_{ij})-\phi(q_{ij},x_{ij})s_{ij}\right)\left(\phi(w,x_{jk})-\phi(q_{jk},x_{jk})s_{jk}\right)=\\ \\ =\phi(z,x_{ij})\phi(w,x_{jk})\uuline{-\phi(z,x_{ij})\phi(q_{jk},x_{jk})s_{jk}}\\ \\ \uwave{-\phi(q_{ij},x_{ij})\phi(w,x_{ik})s_{ij}}+\phi(q_{ij},x_{ij})\phi(q_{jk},x_{ik})s_{ij}s_{jk}\,,\end{array}$ (4.47)
$\begin{array}[]{ccc}R_{ik}^{w}(q_{i}-q_{k})R_{ij}^{z-w}(q_{i}-q_{j})=\left(\phi(w,x_{ik})-\phi(q_{ik},x_{ik})s_{ik}\right)\left(\phi(z-w,x_{ij})-\phi(q_{ij},x_{ij})s_{ij}\right)=\\ \\ =\phi(w,x_{ik})\phi(z-w,x_{ij})\uwave{-\phi(w,x_{ik})\phi(q_{ij},x_{ij})s_{ij}}\\ \\ \underline{-\phi(q_{ik},x_{ik})\phi(z-w,x_{kj})s_{ik}}+\phi(q_{ik},x_{ik})\phi(q_{ij},x_{kj})s_{ik}s_{ij}\,,\end{array}$ (4.48)
@@ 5.1 The Lax pair
[…]
@@
˙$\begin{array}[]{c}\displaystyle{{\dot{\mathcal{L}}}(z)=[{\mathcal{L}}(z),{\mathcal{M}}(z)]}\end{array}$ (5.12)
@@ 5.2 Proof of the Lax equation
~$\begin{array}[]{c}\displaystyle{\tilde{{\mathcal{M}}}^{aa}_{ii}(z):={\mathcal{M}}^{aa}_{ii}(z)+{\mathcal{H}}=A_{i}+{\mathcal{H}}=\frac{1}{2}\sum\limits_{k,l\neq i}^{n}g\Big{(}F^{0}_{kl}(q_{kl})+{\tilde{F}}^{0}_{kl}(q_{kl}^{+})\Big{)}+\frac{1}{2}\sum\limits_{k\neq i}^{n}{\tilde{Y}}_{k}^{0}(q_{k})\,,}\end{array}$ (5.13)
@@
Here we use the notation ~$\tilde{{\mathcal{M}}}_{ii}^{aa}$ from (5.13). The l.h.s. of (5.17) equals $g(\dot{q_{i}}-\dot{q_{j}})F^{z}_{ij}$. For the r.h.s. of (5.17) we have:
@@
Here we have additional terms ${\mathcal{L}}_{ii}^{12}\tilde{\mathcal{M}}_{ii}^{22}-\tilde{\mathcal{M}}_{ii}^{11}{\mathcal{L}}_{ii}^{12}$, which are absent in (3.32). However, the equality ${\mathcal{L}}_{ii}^{12}\tilde{\mathcal{M}}_{ii}^{22}=\tilde{\mathcal{M}}_{ii}^{11}{\mathcal{L}}_{ii}^{12}$ holds true in the operator case as well. This follows from the explicit form of ~$\tilde{\mathcal{M}}_{ii}^{aa}$ given in (5.13) and (4.39), (4.40). Therefore, similarly to (3.33) we have:
@@
~${\tilde{\mathcal{M}}}^{11}(z)={\tilde{\mathcal{M}}}^{22}(-z),\qquad{\mathcal{M}}^{12}(z)={\mathcal{M}}^{21}(-z).$ (5.24)
@@
˙$\dot{{\mathcal{L}}}^{21}(z)={\mathcal{L}}^{21}(z){\tilde{\mathcal{M}}}^{11}(z)+{\mathcal{L}}^{22}(z){\mathcal{M}}^{21}(z)-{\mathcal{M}}^{21}(z){\mathcal{L}}^{11}(z)-{\tilde{\mathcal{M}}}^{22}(z){\mathcal{L}}^{21}(z),$ (5.25)
˙$\dot{{\mathcal{L}}}^{22}(z)={\mathcal{L}}^{21}(z){\mathcal{M}}^{12}(z)+{\mathcal{L}}^{22}(z){\tilde{\mathcal{M}}}^{22}(z)-{\mathcal{M}}^{21}(z){\mathcal{L}}^{12}(z)-{\tilde{\mathcal{M}}}^{22}(z){\mathcal{L}}^{22}(z)$ (5.26)
@@ 6 Representation via the Baxter’s 8-vertex $R$-matrix
$\begin{array}[]{c}R_{12}^{\hbar}(z)=\displaystyle{\frac{1}{2}}\left(\begin{array}[]{cccc}\varphi_{00}+\varphi_{10}&0&0&\varphi_{01}-\varphi_{11}\\ 0&\varphi_{00}-\varphi_{10}&\varphi_{01}+\varphi_{11}&0\\ 0&\varphi_{01}+\varphi_{11}&\varphi_{00}-\varphi_{10}&0\\ \varphi_{01}-\varphi_{11}&0&0&\varphi_{00}+\varphi_{10}\end{array}\right)\,,\end{array}$ (6.3)
@@
where $\nu_{4}=\nu_{0}$ is assumed, and $\partial_{\tau}\omega_{k}$ is equal to either 00 (for $\omega_{0}$, $\omega_{1}$) or $1/2$ (for $\omega_{2}$, $\omega_{3}$). Up to simple redefinitions it coincides with the known $K$-matrix from [12, 17]. It satisfies the property (4.18) that can be deduced from (A.16) and the quasi-periodic behaviour (A.25)-(A.28). Notice also that ${\tilde{K}}^{\hbar}(z)={\tilde{K}}^{z}(\hbar)$, that is ${K}^{\hbar}(z)={\tilde{K}}^{\hbar}(z)$ in this description.
@@ 7 Concluding remarks: XYZ long-range spin chain of ${\rm BC}_{n}$ type
$\begin{array}[]{c}\displaystyle{p_{i}=0\,,\quad q_{i}=z_{i}\,,\ i=1,...,n}\end{array}$ (7.1)
@@
${g^{2}\sum_{k:k\neq i}^{n}\Big{(}\wp^{\prime}(z_{i}-z_{k})+\wp^{\prime}(z_{i}+z_{k})\Big{)}+\frac{1}{2}\sum\limits_{a=0}^{3}\nu_{a}^{2}\wp^{\prime}(z_{i}+\omega_{a})=0\,,\quad i=1,...,n\,.}$ (7.2)
To the best of our knowledge, these equilibrium positions are unknown in the general case. At the same time, it is known that $z_{i}$ coincide with positions of zeros of solutions to the Heun equation (in the elliptic form) [8]. Consider the restriction of ${\mathcal{L}}(z)$ and ${\mathcal{M}}(z)$ (5.1)-(5.11) to the positions (7.1):
${{\mathcal{L}}^{\prime}(z)={\mathcal{L}}(z)|_{p_{j}=0,q_{j}=z_{j}}\,,\quad{\mathcal{M}}^{\prime}(z)={\mathcal{M}}(z)|_{q_{j}=z_{j}}\,,\quad{{\mathcal{H}}^{0}}^{\prime}={\mathcal{H}}^{0}|_{q_{j}=z_{j}}\,.}$ (7.3)
@@
$\begin{array}[]{c}\displaystyle{{{\mathcal{H}}^{0}}^{\prime}=g\sum_{k<l}^{n}\Big{(}F^{0}_{kl}(z_{k}-z_{l})+{\tilde{F}}^{0}_{kl}(z_{k}+z_{l})\Big{)}+\frac{1}{2}\sum\limits_{k=1}^{n}\tilde{Y}_{k}^{0}(z_{k})=}\\ \displaystyle{=\frac{g}{2}\sum_{k<l}^{n}\Big{[}-E_{2}(z_{k}-z_{l})\stackrel{{\scriptstyle k}}{{\sigma}}_{0}\otimes\stackrel{{\scriptstyle l}}{{\sigma}}_{0}+}\\ \displaystyle{+\sum\limits_{a=1}^{3}\varphi_{a}(z_{k}-z_{l},\omega_{a})\Big{(}E_{1}(z_{k}-z_{l}+\omega_{a})-E_{1}(z_{k}-z_{l})-E_{1}(\omega_{a})\Big{)}\stackrel{{\scriptstyle k}}{{\sigma}}_{4-a}\otimes\stackrel{{\scriptstyle l}}{{\sigma}}_{4-a}\Big{]}-}\\ \displaystyle{-\sum\limits_{k=1}^{n}\Big{[}\nu_{0}E_{2}(z_{k})\stackrel{{\scriptstyle k}}{{\sigma}}_{0}+\sum\limits_{a=1}^{3}\nu_{a}\varphi_{a}(z_{k}-\omega_{a},\omega_{a})\Big{(}E_{1}(z_{k}+\omega_{a})-E_{1}(z_{k})-E_{1}(\omega_{a})\Big{)}\stackrel{{\scriptstyle k}}{{\sigma}}_{4-a}\Big{]}\,.}\end{array}$ (7.5)

### v3

@@ 3 Takasaki’s $2n\times 2n$ Lax pair for Calogero-Inozemtsev system
Here we recall the Lax pair for the Calogero-Inozemtsev model (1.11) with five arbitrary constants $g,\nu_{0},\nu_{1},\nu_{2},\nu_{3}\in\mathbb{C}$. Main purpose of this Section is to write down detailed verification of the classical Lax equation for the Takasaki’s $2n\times 2n$ Lax pair. In the next Section this proof will be generalized to $R$-matrix valued case.
@@ 3.1 The model and the Lax pair
Equations of motion are of the form:
@@
In [13] $3n\times 3n$ Lax representation was suggested. Some particular cases (not arbitrary constants) were previously known from [26], see also [4]. Here we deal with $2n\times 2n$ Lax representation proposed by K. Takasaki in [33]. Then O. Chalykh suggested a derivation of quantum and classical Lax pairs from Dunkl operators in [3]. His construction works for the Calogero-Inozemtsev system and for its relativistic generalization as well, so we use the expression from [3]. The description of the spectral curve can be found in [1], and the case of $n=1$ was also studied in [35].
@@
where all entries $L^{ab}(z)$ and $M^{ab}(z)$, $a,b=1,2$ are $n\times n$ matrices (below i,j=1,,ni,j=1,...,n):
@@
where the functions $\phi$ and $f$ are given by (2.5) and (A.7) respectively and the following notation are used:
@@
[…]
@@ 3.2 Lax equation
The explicit form can be obtained from (3.30) by replacing $q_{j},p_{j}$ to $q_{j},-p_{j}$.
@@ 4 Shibukawa–Ueno elliptic $R$-operator and $K$-matrices
In this Section we formulate the parametric version of the Kirillov’s $B$-type associative Yang-Baxter equations. We show that these relations are fulfilled with the operator-valued Shibukawa-Ueno elliptic $R$-matrix and the Komori-Hikami $K$-matrix.
@@ 4.1 Associative Yang-Baxter equation and $K$-operators
The operator $R_{ij}$ acts on a space of meromorphic functions of variables $x_{1},\dots x_{n}$ (instead of a tensor product of vector spaces). Here, the symmetric group $S_{N}$ acts by exchanging the variables $x_{i}$ and $\hat{s}_{ij}$ denotes the corresponding transposition
@@
Introduce another operator $\tilde{K}$:
@@
$\scriptstyle\blacksquare$
Let us introduce the notations $F_{ij}^{z}(q)$ (c.f. (2.8)) and $\tilde{F}_{ij}^{z}(q)$ for the corresponding partial derivatives of $R_{ij}^{z}(q)$ and $\tilde{R}_{ij}^{z}(q)$:
@@
Proof. Taking the partial derivative with respect to $q$ in (4.15), (4.16), (4.17), (4.18), we obtain (4.25), (4.26), (4.27), (4.28). $\scriptstyle\blacksquare$
@@ 4.2 ${\rm BC}_{n}$ associative Yang-Baxter equations with parameters
In order to define the associative classical Yang-Baxter algebra of type ${\rm A}_{n}$ generated by the elements $r_{ij}$ for $1\leq i\neq j\leq n$, we extend (4.29) with the antisymmetry condition and the commutativity relation:
@@
The equation (1.8) with parameters was suggested by A. Polishchuk in [28]. The relation (4.29) looks as half of the classical Yang-Baxter equation. From (4.29) and (4.30) the classical Yang-Baxter equation follows:
$r_{12},r_{13}]+[r_{12},r_{23}]+[r_{13},r_{23}]=0.$ (4.31)
@@
Proof. The proof is straightforward, one should use the definitions (4.3), (4.5) and move all operators $\hat{s}_{ij}$ and $\hat{t}_{i}$ to the left using relations (4.6). The relation (4.32) follows from (A.37). The relations (4.34) follow from the Fay identity (2.4), (A.37) and the equality in the symmetric group $s_{ij}s_{jk}=s_{ik}s_{ij}=s_{jk}s_{ik}$. The four term relation (4.35) follows form (3.22). $\scriptstyle\blacksquare$
@@
[…]
@@ 5 $R$-matrix-valued generalization of Takasaki’s Lax pair
Below we suggest $R$-matrix valued generalization of the Takasaki’s Lax pair (3.2)-(3.10). That is we deal with $2n\times 2n$ Lax pair which entries are operators acting on functions of $n$ variables as in (4.1) and (4.4).
@@ 5.1 The Lax pair
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}(z)=\left(\begin{array}[]{cc}{{\mathcal{L}}^{11}(z)}&{{\mathcal{L}}^{12}(z)}\\ \\ {{\mathcal{L}}^{21}(z)}&{{\mathcal{L}}^{22}(z)}\end{array}\right)\,,\qquad{\mathcal{M}}(z)=\left(\begin{array}[]{cc}{{\mathcal{M}}^{11}(z)}&{{\mathcal{M}}^{12}(z)}\\ \\ {{\mathcal{M}}^{21}(z)}&{{\mathcal{M}}^{22}(z)}\end{array}\right)+{\mathcal{H}}\,1_{2n\times 2n}\,,}\end{array}$ (5.1)
@@
$\begin{array}[]{c}\displaystyle{{\dot{\mathcal{L}}}(z)=[{\mathcal{L}}(z),{\mathcal{M}}(z)]}\end{array}$ (5.12)
@@ 5.2 Proof of the Lax equation
$\begin{array}[]{c}\displaystyle{\tilde{{\mathcal{M}}}^{aa}_{ii}(z):={\mathcal{M}}^{aa}_{ii}(z)+{\mathcal{H}}=A_{i}+{\mathcal{H}}=\frac{1}{2}\sum\limits_{k,l\neq i}^{n}g\Big(F^{0}_{kl}(q_{kl})+{\tilde{F}}^{0}_{kl}(q_{kl}^{+})\Big)+\frac{1}{2}\sum\limits_{k\neq i}^{n}{\tilde{Y}}_{k}^{0}(q_{k})\,,}\end{array}$ (5.13)
@@
Here we use the notation $\tilde{{\mathcal{M}}}_{ii}^{aa}$ from (5.13). The l.h.s. of (5.17) equals $g(\dot{q_{i}}-\dot{q_{j}})F^{z}_{ij}$. For the r.h.s. of (5.17) we have:
@@
Here we have additional terms ${\mathcal{L}}_{ii}^{12}\tilde{\mathcal{M}}_{ii}^{22}-\tilde{\mathcal{M}}_{ii}^{11}{\mathcal{L}}_{ii}^{12}$, which are absent in (3.32). However, the equality ${\mathcal{L}}_{ii}^{12}\tilde{\mathcal{M}}_{ii}^{22}=\tilde{\mathcal{M}}_{ii}^{11}{\mathcal{L}}_{ii}^{12}$ holds true in the operator case as well. This follows from the explicit form of $\tilde{\mathcal{M}}_{ii}^{aa}$ given in (5.13) and (4.39), (4.40). Therefore, similarly to (3.33) we have:
@@
${\tilde{\mathcal{M}}}^{11}(z)={\tilde{\mathcal{M}}}^{22}(-z),\qquad{\mathcal{M}}^{12}(z)={\mathcal{M}}^{21}(-z).$ (5.24)
@@
$\dot{{\mathcal{L}}}^{21}(z)={\mathcal{L}}^{21}(z){\tilde{\mathcal{M}}}^{11}(z)+{\mathcal{L}}^{22}(z){\mathcal{M}}^{21}(z)-{\mathcal{M}}^{21}(z){\mathcal{L}}^{11}(z)-{\tilde{\mathcal{M}}}^{22}(z){\mathcal{L}}^{21}(z),$ (5.25)
$\dot{{\mathcal{L}}}^{22}(z)={\mathcal{L}}^{21}(z){\mathcal{M}}^{12}(z)+{\mathcal{L}}^{22}(z){\tilde{\mathcal{M}}}^{22}(z)-{\mathcal{M}}^{21}(z){\mathcal{L}}^{12}(z)-{\tilde{\mathcal{M}}}^{22}(z){\mathcal{L}}^{22}(z)$ (5.26)
@@ 6 Representation via the Baxter’s 8-vertex $R$-matrix
[…]
@@
where $\nu_{4}=\nu_{0}$ is assumed, and $\partial_{\tau}\omega_{k}$ is equal to either 0 (for $\omega_{0}$, $\omega_{1}$) or $1/2$ (for $\omega_{2}$, $\omega_{3}$). Up to simple redefinitions it coincides with the known $K$-matrix from [12, 17]. It satisfies the property (4.18) that can be deduced from (A.16) and the quasi-periodic behaviour (A.25)-(A.28). Notice also that ${\tilde{K}}^{\hbar}(z)={\tilde{K}}^{z}(\hbar)$, that is ${K}^{\hbar}(z)={\tilde{K}}^{\hbar}(z)$ in this description.
@@ 7 Concluding remarks: XYZ long-range spin chain of ${\rm BC}_{n}$ type
$\begin{array}[]{c}\displaystyle{p_{i}=0\,,\quad q_{i}=\zeta_{i}\,,\ i=1,...,n}\end{array}$ (7.1)
@@
${g^{2}\sum_{k:k\neq i}^{n}\Big(\wp^{\prime}(\zeta_{i}-\zeta_{k})+\wp^{\prime}(\zeta_{i}+\zeta_{k})\Big)+\frac{1}{2}\sum\limits_{a=0}^{3}\nu_{a}^{2}\wp^{\prime}(\zeta_{i}+\omega_{a})=0\,,\quad i=1,...,n\,.}$ (7.2)
To the best of our knowledge, these equilibrium positions are unknown in the general case. At the same time, it is known that $\zeta_{i}$ coincide with positions of zeros of solutions to the Heun equation (in the elliptic form) [8]. Consider the restriction of ${\mathcal{L}}(z)$ and ${\mathcal{M}}(z)$ (5.1)-(5.11) to the positions (7.1):
${{\mathcal{L}}^{\prime}(z)={\mathcal{L}}(z)|_{p_{j}=0,q_{j}=\zeta_{j}}\,,\quad{\mathcal{M}}^{\prime}(z)={\mathcal{M}}(z)|_{q_{j}=\zeta_{j}}\,,\quad{{\mathcal{H}}^{0}}^{\prime}={\mathcal{H}}^{0}|_{q_{j}=\zeta_{j}}\,.}$ (7.3)
@@
$\begin{array}[]{c}\displaystyle{{{\mathcal{H}}^{0}}^{\prime}=g\sum_{k<l}^{n}\Big(F^{0}_{kl}(\zeta_{k}-\zeta_{l})+{\tilde{F}}^{0}_{kl}(\zeta_{k}+\zeta_{l})\Big)+\frac{1}{2}\sum\limits_{k=1}^{n}\tilde{Y}_{k}^{0}(\zeta_{k})=}\\ \displaystyle{=-\frac{g}{2}\sum_{k<l}^{n}\Big(E_{2}(\zeta_{k}-\zeta_{l})+E_{2}(\zeta_{k}+\zeta_{l})\Big)\stackrel{{\scriptstyle k}}{{\sigma}}_{0}\otimes\stackrel{{\scriptstyle l}}{{\sigma}}_{0}+}\\ \displaystyle{+\frac{g}{2}\sum_{k<l}^{n}\sum\limits_{a=1}^{3}\Big[\varphi_{a}(\zeta_{k}-\zeta_{l},\omega_{a})\Big(E_{1}(\zeta_{k}-\zeta_{l}+\omega_{a})-E_{1}(\zeta_{k}-\zeta_{l})-E_{1}(\omega_{a})\Big)+}\\ \displaystyle{+\varphi_{a}(\zeta_{k}+\zeta_{l},\omega_{a})\Big(E_{1}(\zeta_{k}+\zeta_{l}+\omega_{a})-E_{1}(\zeta_{k}+\zeta_{l})-E_{1}(\omega_{a})\Big)\Big]\stackrel{{\scriptstyle k}}{{\sigma}}_{4-a}\otimes\stackrel{{\scriptstyle l}}{{\sigma}}_{4-a}-}\\ \displaystyle{-\frac{1}{2}\sum\limits_{k=1}^{n}\Big[\nu_{0}E_{2}(\zeta_{k})\stackrel{{\scriptstyle k}}{{\sigma}}_{0}+\sum\limits_{a=1}^{3}\nu_{a}\varphi_{a}(\zeta_{k}-\omega_{a},\omega_{a})\Big(E_{1}(\zeta_{k}+\omega_{a})-E_{1}(\zeta_{k})-E_{1}(\omega_{a})\Big)\stackrel{{\scriptstyle k}}{{\sigma}}_{4-a}\Big]\,.}\end{array}$ (7.5)
@@
Main notations:
$n$ – number of particles in the elliptic Calogero-Moser (1.1) and the Calogero-Inozemtsev model (1.11);
$q_{i}$ – the $i$-th particle coordinate;
$p_{i}$ – the $i$-th particle momenta;
$q_{ij}$ and $q_{ij}^{+}$ – short notation for $q_{i}-q_{j}$ and $q_{i}+q_{j}$, respectively;
$\zeta_{i}$ – equilibrium positions, see (2.18) and (7.1);
$\nu_{a}$ for $a=0,1,2,3$ – complex parameters in the Calogero-Inozemtsev model (1.11)
NN – rank of ${\rm GL}_{N}$ $R$-matrix;
$\tau$ – moduli of elliptic curve $\mathbb{C}/\Gamma$, ${\rm Im}(\tau)>0$;
$\phi(z,u)$ – Kronecker elliptic function $\phi(z,u)$ (1.6), (A.1);
$\wp(z)$ – the Weiestrass elliptic $\wp$-function;
$v(z,u)$ –elliptic function (3.19) depending on the parameters $\nu_{0},\nu_{1},\nu_{2},\nu_{3}$ ;
$\hbar$ – complex quantum parameter entering the $R$-operator definition;
$z$ – spectral parameter in Lax matrix, at the same time the quantum parameter in $R$-operators in $R$-matrix Lax pairs;
$R^{\hbar}(u)$ — elliptic $R$-operator, satisfying (1.8), (1.9), (2.5) and (2.6). Depending on the context, it is the Shibukawa-Ueno operator (4.1), which acts on the space of scalar functions on $n$-variables, or its finite-dimensional analogue, the Baxter-Belavin elliptic $R$-matrix, acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{N}$, see (6.1) for $N=2$;
$_{ij}^{\hbar}(u)RR$-operator, acting nontrivially on variables $x_{i}$ and $x_{j}$, or $R$-matrix, i.e. an element of ${\rm End}((\mathbb{C}^{N})^{\otimes n})$ acting on the $i$-th and the $j$-th tensor components of $(\mathbb{C}^{N})^{\otimes n}$;
$\tilde{K}_{i}^{\hbar}(u)$ – the corresponding $K$-operator, satisfying (1.15) and (4.14), see (4.12) and (6.5);
$F_{ij}^{\hbar}(u)$ – partial derivative of $R$-operator $\partial_{u}R_{ij}^{\hbar}(u)$;
$\tilde{Y}_{i}^{\hbar}(u)$ – partial derivative of $K$-operator $\partial_{u}K_{i}^{\hbar}(u)$;
$R_{ij}^{\hbar}$, $F_{ij}^{\hbar}$ –short notation for $R_{ij}^{\hbar}(q_{i}-q_{j})$ and $F_{ij}^{\hbar}(q_{i}-q_{j})$;
$\tilde{K}_{ij}^{\hbar}$, $\tilde{Y}_{ij}^{\hbar}$ –short notation for $\tilde{K}_{i}^{\hbar}(q_{i})$ and $\tilde{Y}_{i}^{\hbar}(q_{i})$;
$L(z)$ and $M(z)$ – Lax pair in the scalar case, see (1.5) and (2.1) for the elliptic Calogero-Moser and (3.2) for the Calogero-Inozemtsev model;
${\mathcal{L}}(z)$ and $\bar{\mathcal{M}}(z)$$R$-matrix valued Lax pair, see (2.12)-(2.14) for the elliptic Calogero-Moser and (5.1) for the Calogero-Inozemtsev model.